OpenStudy (anonymous):

someone tell me what i did wrong divide and simplify z^2-49 divided by 2z-14 ------- ------ z^2-8z+16 z^2-2z-63 z^2-49 z^2-2z-63 ------ * ---------- z^2-8z+16 2z-14 (z^2-49)(z^2-2z-63) ------------------- (z^2-8z+16)(2z-14) this is were im stuck

OpenStudy (anonymous):

note that 49 = 7^2 so z^2 -49 = z^2-7^2 also 2z-14 = 2(z-7) also, you should factorize z^2-8z+16 and z^2-2z-63

OpenStudy (anonymous):

so what do i do next

OpenStudy (anonymous):

factorize them. do you know how to factorize?

OpenStudy (anonymous):

(z-7)(z+7)(z-7)(z+1) ---------------- (z-4)(z+4)(2)(z-7)

OpenStudy (anonymous):

no, not entirely correct. try again.

OpenStudy (anonymous):

(z-7)(z+7)(z-1)(z+2)(z-7)(z+9) ------------------------ (z-4)(z+4)(2)(z-7)

OpenStudy (anonymous):

how did you get 6 terms in your numerator?

OpenStudy (anonymous):

-7,7=-49 -1,2=-2 -7,9=-63

OpenStudy (anonymous):

yes, but why is the -1,2 = 2 there?

OpenStudy (anonymous):

I don't see any need for that.

OpenStudy (anonymous):

i thought it was suppose to be when i took it out you said it was wrong

OpenStudy (anonymous):

no, I said your factorization was incorrect.

OpenStudy (anonymous):

(z^2-49) = (z+7)(z-7) (z^2-2z-63) = (z-9)(z+7) (z^2-8z+16) = (z-4)^2 2z-14 = 2(z-7) I hope you can take it from there.

OpenStudy (anonymous):

then rewrite and simplify (z-7)(z+7)(z-9)(z-7) ---------------- (z-7)(z-4)(2)(z-7) (z-7) ---- 1*(z+7)(z-9)(z+7) (z-7) ---------- (z-4)(2)(z-7)